A rigid tank contains 54.5 g of chlorine gas (Cl2) at a temperature of 72 °C and an absolute pressure of 5.60×10^5 Pa. The mass per mole of Cl2 is 70.9 g/mol. a) Determine the volume of the tank. b) Later, the temperature of the tank has dropped to 33 °C and, due to a leak, the pressure has dropped to 3.40×10^5 Pa. How many grams of chlorine gas has leaked out of the tank? (The mass per mole of Cl2 is 70.9 g/mol.)8. A rigid tank contains 54.5 g of chlorine gas (Cl2) at a temperature of 72 °C and an absolute pressure of 5.60 x 10 Pa. Th
Answer #1

massaf ells, M=.54. 59 temp. Ti = 72+273 = 345k Pressur P,= 5.60 x10 Spa Malan massages, m=70.906 almal 1.470.9 g/mol @ idea

b final temp, Is = 33° +273 = 300k P2= 3.foxlo Spa V=V2 let ho. of mates escaped be n. P2 Y = (54.5 - 2)x 8.31 4X Tz 3- 40x10